OpenStudy (anonymous):

Find three consecutive integers such that the first increased by twice the second is 24 more than the third.

OpenStudy (turingtest):

first off you will need a way to represent three consecutive integers with one variable any idea how to do that?

OpenStudy (anonymous):

We are normally good with these problems, but the wording messed us up. So, we could find the numbers if someone could help us find out what we need to do.

OpenStudy (turingtest):

Like I say, the fist step is to say "how can I represent this without 3 variables, because we are talking about 3 different numbers" But they are consecutive, so we can write them in a special way. n n+1 n+2 that is how we will talk about our first number, second number and third number...

OpenStudy (anonymous):

Hmmm.... We are going to try the problem again, because we got a non- integer too.

myininaya (myininaya):

ok so we have three consecutive integers right

myininaya (myininaya):

just like turing pointed out above n,n+1,n+2

OpenStudy (anonymous):

Yes, we just don't understand why it said that the second is 24 more than the third. That lost us.

myininaya (myininaya):

is means equals

myininaya (myininaya):

24 more than the third means 24+third integer

myininaya (myininaya):

the third integer here is n+2

myininaya (myininaya):

so the second part of the equation is =24+(n+2)

OpenStudy (anonymous):

Okay that makes more sense.

myininaya (myininaya):

so you have the first part of the equation?

myininaya (myininaya):

can you tell me what it is?

OpenStudy (anonymous):

So, we are going to take a guess right now because one of us is writing this out while one of us is typing. Okay, So, since the second part is = 24 (n+2) then would that mean that the whole thing is 2n(n+1)=24(n+2)?

myininaya (myininaya):

oh no

myininaya (myininaya):

the second part is =24+(n+2)

myininaya (myininaya):

not =24(n+2)

OpenStudy (anonymous):

Okay, Sorry we forgot the + ^^;;;

myininaya (myininaya):

the first part of the equation is first integer + 2(second integer)

OpenStudy (anonymous):

So, umm.. n+2(n+1) = 24+(n+2)?

myininaya (myininaya):

yep yep

OpenStudy (anonymous):

Oh! Okay! 8D

myininaya (myininaya):

So you are good at solving these?

OpenStudy (anonymous):

We normally can, we just had a issue with the figuring of the equation. The wording was odd to us. We could solve it and then post it?

myininaya (myininaya):

ok awesome

OpenStudy (anonymous):

n+2(n+1)=24+(n+2) n+2n+2= 24+n+26 3n+2=50+n -n -n ________________ 2n+2=50 -2 -2 ________________ 2n=48 /2 /2 ________________ n= 24

myininaya (myininaya):

one sec while i check

myininaya (myininaya):

n+2(n+1)=24+(n+2) n+2n+2= 24+n+26 <---- how did you get this 26 3n+2=50+n -n -n ________________ 2n+2=50 -2 -2 ________________ 2n=48 /2 /2 ________________ n= 24

OpenStudy (anonymous):

24+2 We added there.

myininaya (myininaya):

oh

myininaya (myininaya):

lol

myininaya (myininaya):

wait

myininaya (myininaya):

why did you leave 24 there?

myininaya (myininaya):

after adding the 2

OpenStudy (anonymous):

24+(n+2) 24+n for one And 26 for the other So you have 24+n+26

OpenStudy (turingtest):

No you do not "distribute" with addition, only with multiplication.

myininaya (myininaya):

so we have 24+(n+2) is the same as 24+n+2 24+2+n

OpenStudy (anonymous):

So, we should have done the second one we did. n+2(n+1)=24+(n+2) n+2n+2= 24+n+2 3n+2=26+n -n -n ________________ 2n+2=26 -2 -2 ________________ 2n=24 /2 /2 ________________ n= 12

myininaya (myininaya):

yes great! :)

OpenStudy (anonymous):

We mean we should have gone with.

OpenStudy (anonymous):

^^! Thank you very much for helping us out!

myininaya (myininaya):

no problem thanks for wanting to learn and not just wanting the answer thats really cool

myininaya (myininaya):

I just became your fan. :)

OpenStudy (anonymous):

No problem. We are always interested in the way things work. :) We'll return the favor! And if you ever need help in things like History and English those are our best subjects. Hopefully we can get around to helping more since we just got on. ^^ Thank you again!

myininaya (myininaya):

Great! Happy Holidays!

OpenStudy (anonymous):

You too! And a Happy New Year!

myininaya (myininaya):

You too! lol